OpenStudy (anonymous):

how do i wite a formula for this $12000 invested @ 6% compounded semi annually fo 5 years thanks in advance

OpenStudy (cwrw238):

Amount = 12,000( 1 + 0.06/2)^(2*5) = 12000(1 + 0.03)^10

OpenStudy (anonymous):

ty again

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

investmant percents*semi then what is the 2*5 pls

OpenStudy (cwrw238):

the general form is A = P(1 + r/c )^cr where P = principal amount, r = rate per annum and c is the number of times interest is paid per annum

OpenStudy (cwrw238):

the 2 in 2*5 is the number of times its paid per annum

OpenStudy (anonymous):

crw if you are still there i would like to say a big thank you, my laptop started a load of pop ups, throwed it at the wall and you can guess what happened so now using a spare,

OpenStudy (cwrw238):

oh ok you are welcome

OpenStudy (anonymous):

ty just got so frustrated with it its been doing that for past 2 days finally snapped lol good job it was not brand new

OpenStudy (anonymous):

A = P(1 + r/c )^cr 12000(1+0.06/2)^? sorry just cant seem to get the ^cr yet to get to my answer

OpenStudy (anonymous):

2*5

OpenStudy (anonymous):

got it i think

OpenStudy (anonymous):

bye